Calculate the pH of 0.0800 mol dm ^-3 C2H5COOH(aq)
1
Expert's answer
2020-05-14T14:33:10-0400

C2H5COOH, or propionic acid, dissociates in water solution according to the equation:

C2H5COOH + H2O "\Leftrightarrow" H3O+ + C2H5COO-.

The acid dissociation constant expression involves the equilibrium concentrations of protonated, deprotonated forms and of hydronium cation:

"K_a = \frac{[H_3O^+][C_2H_5COO^-]}{[C_2H_5COOH]}" .

The Ka of propionic acid is 10-4.88. If the initial concentration of the acid was 0.0800 M and the equilibrium concentration of H3O+ and C2H5COO- are "x", the Ka is:

"K_a = \frac{x^2}{0.0800 - x}".

Then , the equilibrium concentration of hydronium ion is:

"x^2 + K_a\u00b7x - 0.0800\u00b7K_a = 0"

"x = 0.00102"

"[H_3O^+] = 0.00102" M.

Finally, the pH of the solution is:

"pH = -\text{log}[H_3O^+]"

"pH = 2.99".

Answer: the pH of 0.0800 mol dm-3 C2H5COOH(aq) is 2.99.